### v2

# Sparse Feature Coactivation Reveals Causal Semantic Modules in Large Language Models

Abstract
We identify semantically coherent, context-consistent network components in large language models (LLMs) using coactivation of sparse autoencoder (SAE) features collected from just a handful of prompts. Focusing on concept-relation prediction tasks, we show that ablating these components for concepts (e.g., countries and words) and relations (e.g., capital city and translation language) changes model outputs in predictable ways, while amplifying these components induces counterfactual responses. Notably, composing relation and concept components yields compound counterfactual outputs. Further analysis reveals that while most concept components emerge from the very first layer, more abstract relation components are concentrated in later layers. Lastly, we show that extracted components more comprehensively capture concepts and relations than individual features while maintaining specificity. Overall, our findings suggest a modular organization of knowledge accessed through compositional operations, and advance methods for efficient, targeted LLM manipulation.
Sparse Feature Coactivation Reveals Causal Semantic Modules
in Large Language Models
Ruixuan Deng11thanks: Indicates equal contribution.    Xiaoyang Hu2211footnotemark: 1    Miles Gilberti3311footnotemark: 1    Shane Storks3311footnotemark: 1 Aman Taxali33    Mike Angstadt33    Chandra Sripada33    Joyce Chai33 1Georgia Institute of Technology   2Brown University   3University of Michigan rdeng62@gatech.edu,    xiaoyang_hu@brown.edu, {milgil, sstorks, ataxali, mangstad, sripada, chaijy}@umich.edu
Figure 1: (A) We construct inter-layer feature networks from SAE coactivation patterns, prune high-density features, and extract task-relevant components. (B) Components are often consistent across contexts. (C) Selective ablation and amplification of components steers the model toward counterfactual outputs, overriding the original prompt.
1 Introduction
Contemporary neural large language models (LLMs) exhibit remarkable proficiency in representing and reasoning over relational knowledge, motivating efforts toward mechanistic interpretation. Surprisingly, despite the immense complexity of LLMs, prior work has modeled this capability with simple linear transformations (Hernandez et al., 2024) and vector arithmetic (Merullo et al., 2024). However, these accounts are limited to layerwise interpretations, lacking fine-grained interpretation of specific features and their interaction.
Sparse autoencoders (SAEs; Huben et al., 2024) recently emerged as a powerful tool for extracting interpretable, often monosemantic features from the internal representations of LLMs. SAEs achieve this by learning to reconstruct dense, entangled activations at each neural network layer through sparse codes, where individual features are incentivized to activate rarely and selectively. Despite this progress, it remains unclear how features from different layers organize and coordinate to produce responses (e.g., how an LLM may access and compose its encoded knowledge to respond “Beijing” to a prompt for China’s capital city). Cross-layer transcoders extend SAEs to approximate layerwise computations and globally interpret information flow in LLMs as computation graphs (Dunefsky et al., 2024; Ameisen et al., 2025), these graphs consist of hundreds of densely connected nodes, requiring manual feature grouping for interpretability. As such, there remains a need for a global, sparse, feature-level interpretation of LLMs’ relational knowledge and reasoning.
In this work, we employ SAE feature coactivation to construct inter-layer feature networks for individual prompts about various factual and linguistic relational reasoning tasks. By filtering out denser features that activate across diverse contexts, we identify semantically coherent connected components within these networks representing specific concepts and relations across different prompts. We show that selective ablation and amplification of these components changes model outputs in predictable ways, overriding in-prompt instructions. In particular, composing a relation component with a concept component produces compound counterfactual outputs. As exemplified in Figure 1, a prompt about China’s capital activates components related to China and capital cities, eliciting the correct response of “Beijing” from the model. However, ablating the China component while amplifying the Nigeria component effectively steers its response to “Abuja”, whereas ablating the capital component while amplifying the language component steers it to “Chinese.” Furthermore, simultaneously ablating the China and capital components while amplifying the Nigeria and language components changes the model output to “English.”
In addition to contributing a sparse, modular structure that causally accounts for LLM relational reasoning processes, these components readily enable fine-grained interpretation. For example, they reveal the prominence and topology of specific concepts and relations; by ablating individual component features, we discover that country components tend to emerge from the very first network layer. In contrast, more abstract components for facts about countries are concentrated in later layers. Even within individual country fact components, we observe that nodes from deeper layers tend to have a stronger causal impact on model outputs. This aligns with our understanding of artificial neural networks and biological systems, where higher-level abstractions emerge in later processing stages. Lastly, our analysis demosntrates that our components offer a more comprehensive account of LLM relational reasoning than SAE features alone, and that components for each concept-relation pair are fairly specific to the functions we associate them with. Together, these results offer insights into the organization of relational knowledge in LLMs and establish sparse feature coactivation as an effective, lightweight approach for mechanistic interpretability and targeted model manipulation.
2 Methods
We focus our analyses on Gemma 2 2B (Team et al., 2024) [^1] Accessed via google/gemma-2-2b in Hugging Face Hub (Wolf et al., 2020). LLM activations collected with 1 NVIDIA A100 GPU (40GB VRAM) and 12GB RAM., with some results for Gemma 2 9B in Appendix A. As shown in Figure 1, we perform several steps to construct an inter-layer feature network based on coactivation patterns, extract task-relevant components, and perform targeted causal interventions to assess their functional roles.
Activation collection.
We run each input prompt through the model integrated with pre-trained SAEs [^2] We use SAE Lens to apply the width_16k/canonical variant from the gemma-scope-2b-pt-res-canonical release (Lieberum et al., 2024; Bloom et al., 2024). Each SAE maps the residual stream activation at layer $\ell$, $x_{\ell}\in\mathbb{R}^{d_{\text{model}}}$, to a sparse representation $\phi_{\ell}\in\mathbb{R}^{d_{\text{sae}}}$ where $d_{\text{sae}}=16384$. This produces an activation tensor $\Phi_{\ell}\in\mathbb{R}^{T\times d_{\text{sae}}}$ for each layer $\ell$, where $T$ is the number of non-BOS tokens in the prompt.
Feature selection.
To ensure computational tractability while preserving key information, we select a set $S_{\ell}$ of top-activated features for each layer. A feature index $i$ is included in $S_{\ell}$ if it appears in the top $k=5$ activations at any token position $t\in\{1,\dots,T\}$:
$S_{\ell}=\bigcup_{t=1}^{T}\{i\mid\Phi_{\ell}[t,i]\in\text{top-}k(\Phi_{\ell}[t,:])\}$
Graph construction.
We construct a directed graph $G=(V,E)$ where each node $(\ell,i)\in V$ corresponds to a selected feature $i\in S_{\ell}$. Edges $E$ connect nodes in adjacent layers according to the temporal correlation of their activation patterns across tokens in the prompt. Specifically, for feature pairs in adjacent layers $i\in S_{\ell}$ and $j\in S_{\ell+1}$, we compute the Pearson correlation coefficient $\rho$. A directed edge $e=((\ell,i),(\ell+1,j))$ is added to $E$ if $\rho(\Phi_{\ell}[:,i],\Phi_{\ell+1}[:,j])>\tau_{\text{corr}}=0.9$. Edge weights are assigned as $w(e)=\min(1.0,\rho)$.
Pruning and component identification.
Some SAE features activate frequently across unrelated contexts, making them overly generic and hard to interpret. To reduce noise, we prune the graph using activation density scores from Neuronpedia (Lin, 2023). For each node $(\ell,i)\in V$, we retrieve its activation density $d_{\ell,i}$, defined as the fraction of tokens in a large corpus where the feature activates. We retain only sparse features, i.e., those with $d_{\ell,i}\leq\tau_{\text{density}}=0.01$, [^3] This follows Neuronpedia’s standard, which classifies features below this density as sparse and interpretable. creating a pruned graph $G_{\text{sparse}}$. We also remove any isolated nodes from $G_{\text{sparse}}$. We then use a straightforward BFS-based method implemented by NetworkX [^4] https://networkx.org/ to identify weakly connected components within $G_{\text{sparse}}$.
Causal validation.
To evaluate the functional significance of each component, we perform targeted interventions using TransformerLens (Nanda and Bloom, 2022). Specifically, we ablate or amplify the activations of SAE features in a given component during the model’s forward pass and measure the resulting shift in the probability distribution over next-token predictions. We quantify this shift using KL divergence between the original and perturbed distributions. A component is considered causal if its manipulation leads to systematic changes in model behavior, measured in Section 3.
Figure 2: We extract components from LLM queries to translate love into Spanish, French, and German. Each component is plotted with its feature count on the $x$-axis and the KL divergence between pre- and post-ablation output token distributions on the $y$-axis. Notably, for love and other concept, only a small number of components exert significant causal effects. For love, the top three components for each relation consistently correspond to either the queried word or language.
3 Experimental Results
For our initial experiments, we consider 4 multi-relational prediction tasks where the base model achieves 100% accuracy: (1) capital city, currency, and language of 10 countries, (2) translation of 11 words from English into Spanish, French, and German, and (3) synonym, antonym, past tense, capitalization, and first letter of 8 English verbs. [^5] Latter 2 tasks adapted in part from data used in Todd et al. (2024). See Appendix B for full specifications of tasks.
To collect model activations, we used simple prompt templates with in-context examples, e.g., for each country-capital pair, we used: “The capital city of Peru is Lima. The capital city of South Korea is Seoul. The capital city of Saudi Arabia is Riyadh. The capital city of {country} is”. Similar templates were used across tasks (Appendix C).
The remainder of this section is organized as follows. We first identify semantic components corresponding to each concept and relation (Section 3.1). We elicit counterfactual model outputs by ablating and amplifying these components individually and in combination (Section 3.2). Lastly, we perform several analyses to examine the organization and impact of component features across model layers (Section 3.3), whether our approach enables more reliable manipulation of concepts and relations than individual SAE features (Section 3.4), and whether ablating components for one task degrades model performance on other tasks (Section 3.5).
Capital Spanish Capitalize
Original Ctry. Abl. Fact Abl. Original Word Abl. Lang. Abl. Original Verb Abl. Trans. Abl.
Beijing 97. Madrid 39. the 12. Amor 41. " 5.9 Love 42. Break 22. Capital 10. break 20.
Be .38 Warsaw 10. China 6.7 amor 19. El 4.9 love 14. break 13. The 7.6 Break 7.1
Peking .35 Rome 9.5 Beijing 6.4 Amo 7.0 La 4.4 I 13. BREAK 8.3 4.2 5.1
Abuja 85. New 7.7 Lagos 17. rojo 38. La 4.1 Red 29. Like 21. Capital 5.6 like 12.
Lagos 11. Islamabad 7.3 the 12. Rojo 34. " 3.5 red 16. like 12. The 4.7 6.1
.38 Kathmandu 6.0 Nigeria 7.1 rojo 2.1 Bol 3.2 The 6.4 Capital 8.6 Yes 4.5 capital 5.6
Table 1: Top 3 output tokens and their likelihoods for various concept-relation pairs, before and after ablating concept or relation components. Underscore prefixes of tokens omitted for space. More examples in Appendix H.1.
3.1 Component Identification
For each concept-relation pair, we obtained sparse components using the methods from Section 2. As shown in Figure 2 for translation, [^6] Additional examples in Appendix D. 2-3 of the components typically exerted a markedly higher causal effect on the model output than the others.
Semantic coherence.
To identify semantically coherent representative components for each concept and relation from these most impactful components, we applied a few approaches. First, we inspected their associated feature descriptions from Neuronpedia. [^7] These descriptions from Neuronpedia were generated by prompting GPT-4o mini (OpenAI et al., 2024) to summarize the texts that activate each feature over a large corpus. In many cases, the features within components had thematically coherent descriptions referring to common concepts or relations, e.g., most country fact task components (Appendix E).
However, these descriptions were often unreliable. For instance, none of the components obtained for prompts about Spain mention Spain in their feature descriptions, and some components for the word beautiful included features about Scotland. Additionally, some components for translation languages were described to be about programming. This suggests that even SAE features suffer from a degree of polysemanticity, possibly due to unavoidable limitations in the corpora and batching strategies used for training SAEs. Notably, this shows that it is impossible to consistently extract comparable components by simply performing a keyword search over SAE feature descriptions.
To address this in the country facts task, we performed component ablations and observed the resulting changes in the model’s top predicted tokens, selecting components that yielded the most compositional changes. As an example, the left 3 columns of Table 1 present results from ablation experiments using selected components obtained from prompts for the capital of China and Nigeria. Promisingly, when country components were ablated, the model’s top predicted tokens shifted predictably to the capitals of other countries. When fact components were ablated, the model assigned higher probabilities to country names.
For the remaining tasks, this approach was unsuccessful, possibly due to more complex composition operations at play than can be captured by a single component for a concept or relation. [^8] For example, steering with a particular component for here caused the model to output common locations for the prompted word, e.g., a word meaning ocean for the prompted word fish in both Spanish and French, a nontrivial operation. In these tasks, we instead represented each concept and relation by the union of multiple seemingly relevant causally impactful components. As shown in the other columns of Table 1, ablating these combined components for words/verbs in these tasks similarly caused the model to output other words in the prompted language of Spanish, or other capitalized words. Meanwhile, ablating relation components caused the model to repeat the input word.
Figure 3: Components representing China (blue) and country language (green), visualized by network layer.
Context consistency.
Most concept and relation components were remarkably consistent across different contexts, e.g., the top China components obtained from capital and currency prompts were identical, and the language components from China and Nigeria prompts differed only by a few edges. Therefore, for country facts and translation, we define each concept component as the intersection of all the components for that concept across relations; similarly, each relation component is defined as the intersection of all the components for that relation across concepts. For verb transformations, while many components appeared context-insensitive, using the specific components identified for each prompted concept-relation pair maximized steering performance. Figure 3 visualizes the resulting China and language components, and more examples are provided in Appendix F.
Country Facts Word Translation Verb Transformation
C R Cap. Curr. Lang. Avg. Spanish French German Avg. Syn. Ant. Past T. Cap. 1st L. Avg.
.98 .93 .96 .96 .75 .78 .72 .75 .61 .45 .61 .18 .57 .48
1.0 .90 .90 .93 .91 .95 .91 .92 .00 .00 .00 .38 .91 .26
1.0 .83 .93 .92 .52 .65 .60 .59 .02 .00 .00 .18 .88 .22
Table 2: Steering success rates (SR) for all tasks by relation and averaged. Row 1: SR of steering from one concept to another (C✓) while preserving relation (R✗). Row 2: SR of steering to each relation from others (R✓) while preserving concept (C✗). Row 3: SR of steering to concept-relation pairs from others by target relation (C✓, R✓).
3.2 Component Steering
Having identified distinct components for each concept and relation, we next investigate whether they can be used to steer model outputs individually and in combination. [^9] See Appendix G for details on steering strength selection. To test whether components generalize across different contexts, we applied a zero-shot prompt template different from the one used to collect the initial activations, e.g., for translation: “Q: What is the {Spanish/French/German} word for ’{word}’? Answer directly (two words max). A:”. We used comparable prompts for other tasks (Appendix C). Table 2 summarizes the results by task and relation, while Appendix I further breaks down performance by concepts.
Language, CN French, love Past Tense, break
$\cdot$, NG Capital, $\cdot$ Currency, $\cdot$ $\cdot$, red Spanish, $\cdot$ German, $\cdot$ $\cdot$, like Capitalize, $\cdot$ 1st Letter, $\cdot$
_English 71. _Beijing 96. _Yuan 15. _rouge 44. _amor 41. _Liebe 60. _liked 43. _Past 17. _B 24.
_Yoruba 5.6 _Be 1.5 _Ren 10. _Rouge 30. _el 9.5 _die 4.0 _Liked 43. _Has 7.8 _( 8.2
_Ha 4.8 _Peking .83 _RMB 6.8 _Sang 2.5 _Amor 8.5 _„ 3.3 _L 1.6 _Kick 4.4 _Past 5.8
Table 3: Top 5 output tokens and likelihoods for prompts about China’s language (columns 1–3), love’s French translation (columns 4–6), and break’s past tense (columns 7–9), after ablating an in-prompt concept or relation component (row 1 headings) and amplifying a target concept or relation component (row 2 headings). Highlighted tokens are or may begin correct answers for steered components. More examples in Appendix H.2.
Concept steering.
By ablating an in-prompt concept component and amplifying a target concept component, we successfully directed the LLM to respond to questions about relations for the in-prompt concept with counterfactual answers (i.e., from applying the relation to the target concept which is not actually queried in the prompt) 48-96% of the time (Table 2). For example, as shown in Table 3, the model consistently disregarded prompts’ references to China, love, and break, instead outputting counterfactual answers “English” (Nigeria’s language), “rouge” (red in French), and “liked” (like’s past tense). Even when incorrect answers were among top tokens, they were often relevant to the in-prompt relation and target concept, e.g., the model ranked “Yoruba” and “Ha[[usa]],” (other commonly spoken languages in Nigeria) just behind the correct answer of “English,” while it ranked a French word for blood just behind words for the target concept of red. This confirms that our identified concept components encode concept-specific information that causally determines model outputs.
We observed the most failures in the verb transformations task. Notably, some verbs had much lower concept steering success rate than others, e.g., possess and sink, which had respective concept steering success rates of about 6% and 20%. We observe that these verbs are also highly represented as in-prompt verbs in failures to steer to other target verbs. This suggests that our method did not yield stable components for these particular verbs, causing both failures to ablate and steer based on them. Interestingly, we see when steering to sink from break and hide, the model respectively generates sought and slid, possibly suggesting that components for sink may instead represent words that begin with s. Meanwhile, the capitalize relation had the lowest success rate; we observed that the model usually output the target word, but not capitalized. For hide, the model instead output the target verb in all capital letters. This may suggest that components for specific words can override prompt instructions when used for steering.
Relation steering.
By ablating an in-prompt relation while amplifying a target relation, we also successfully directed the model to respond to queries about the in-prompt relation as though they concerned the target relation 26-93% of the time (Table 2). As shown in Table 3, the model disregarded prompts’ references to country language, French translation, and past tense transformation, and produced correct counterfactual answers for other relations within each task, e.g., answering prompts to translate love into French with Spanish and German words for love. Even when incorrect answers were among top tokens, they were often relevant, e.g., other common languages or words in the target language. The only failure was the target transformation of capitalize, which yielded capitalized words other than the in-prompt verb break.
Unsurprisingly, the most failures again occurred in this task, specifically for the synonym, antonym, and past tense transformations. For the former two, this may be expected. Unlike the capital city or currency of a country, synonym and antonym do not have objective answers, and actually depend on a variety of finer-grained dimensions, e.g., word sense and part of speech. As observed in previous visualized token distributions after ablation and steering, the country fact and word translation components have clear roles in model outputs: to promote tokens related to specific countries or words, or that are relevant to a target fact (languages, currencies, or cities) or in a target language. However, a mechanism for synonym or antonym relations should not be expected to work like this, as many tokens could be a synonym or antonym, and reasonable answers thus depend more heavily on context. While past tense forms of verbs are more consistent, this relation is much broader and less specific than relations in other tasks, as every verb has a past tense form, and thus an LLM having a component to promote past tense inflections may be inefficient. Such relations are likely better captured in attention heads, which aligns with Todd et al. (2024)’s finding that antonym and past tense are modeled well by function vectors.
Currency, China German, love Antonym, break
Capital, Nigeria Language, Nigeria Spanish, red French, red Synonym, like Capitalize, like
_Abuja 70. _English 93. _rojo 45. _rouge 31. _love 46. _Opposite 9.3
_Lagos 26. _French 1.7 _el 9.8 _le 11. _loved 6.2 _Dis 7.9
_ 2.3 _Yoruba 1.5 _roja 4.2 _la 4.9 _lot 6.0 _ 6.3
Table 4: Top 5 output tokens and their likelihoods for prompts about China’s currency (columns 1–2), love’s German translation (columns 3–4), and break’s antonym (columns 5–6), after ablating both in-prompt components and amplifying components for a target concept-relation pair. More examples in Appendix H.3.
Composite steering.
We conducted composite steering experiments where both concept and relation components were manipulated at once. By ablating both the in-prompt concept and in-prompt relation components while amplifying the target concept and target relation components, it was possible to steer the model to ignore both the in-prompt concept and in-prompt relation and answer about a different concept-relation pair 22-92% of the time (Table 2). Table 4 provides specific examples of composite steering success. As shown in the first 2 columns, when we ablated both the China and capital components while amplifying the Nigeria and currency components, the model answered “Naira” despite being asked about China’s capital. In the middle 2 columns, we see similar success in steering the model’s outputs to Spanish and French words for red, despite being prompted for the German translation of love. In the last 2 columns, when prompted for an antonym of break, we successfully steer the model to output a synonym of like, while steering it to output like capitalized (i.e., “Like”) was less successful, with the correct answer only ranked fourth among possible next tokens. Nonetheless, these findings demonstrate the significant composability of our components.
3.3 Component Organization
Having established the causal role and composability of concept and relation components, we next analyze their distribution across network layers and the relative importance of nodes within these components. To quantify the causal importance of an individual concept node, we compute the average post-ablation KL divergence from the original output distribution across all relations. For a relation node, we compute the same average across all concepts. Figure 4 displays results for the country facts task, with more examples in Appendix J. Concept and relation components show distinct distribution patterns across network layers. 8 of 10 country components begin in the first layer, as do all word and verb components in other tasks. Some concept components span nearly all layers (e.g., for China and love), while others concentrate in early to middle layers (e.g., for Nigeria and red). In contrast, relation components concentrate in later layers, with most spanning only the last quarter to half of the network’s layers (e.g., for country facts and target translation languages), and a few spanning across the network (e.g., for verb transformations). This suggests that representations of more concrete entities are established earlier in processing, while more abstract relations emerge later. While concept component nodes exhibit various trends in KL divergence across layers, we find that relation components for the country facts task have stronger causal impact in later layers, while the opposite trend holds for word translation.
Figure 4: KL divergence between pre- and post-ablation output token distributions for each node in the China, Nigeria, and country fact components, plotted by layer. Linear regression lines plotted in red.
3.4 Single Feature Steering Comparison
We hypothesize that our approach of steering based on a group of features is advantageous over steering based on an individual feature (as done in some prior works, e.g., Anthropic, 2023; Templeton et al., 2024). To validate this, we compared its efficacy against a simpler baseline: intervening on only the single most causally impactful feature. Specifically, we identified this top feature for each concept and relation as the one whose individual ablation produced the highest KL divergence, then in line with earlier results, used it for ablation and steering. For the country facts task, the baseline respectively achieved average success rates of 83%, 83%, and 75% for composite country-relation steering. These results fall notably short of the success rates achieved using full components (respectively 96%, 93%, and 92%). This suggests that concepts are represented distributively across populations of features, and our method’s success comes from identifying and manipulating these entire functional groups. Neuroscience research has similarly found evidence for populations of firing neurons encoding information in human brains Xie et al. (2024).
3.5 Cross-Task Evaluation
Lastly, we explore whether ablating our identified components impacts performance on other tasks, an important question for the practicality of manipulating models based on them. Table 5 presents the model’s average accuracy for each task when our identified components are ablated. When ablating components from the country facts task, 100% accuracy (or near it) is maintained in most cases, even when ablating components for some country-fact pairs and prompting the model for other ones. Accuracy on translation degrades slightly, possibly due to the intersection of some country- and language-specific information. In ablating word translation components, accuracy remains high on other tasks, but degrades slightly on translation itself. This may be expected, given that the relations in this task are of highly similar nature, and even overlap by a small number of features. In ablating verb transformation components, we see moderately degraded accuracy for all tasks to as low as 64-73%. This suggests that the components for this task are less specific than other tasks, in line with our relatively low steering performance with them. Nonetheless, this demonstrates that extracted components are generally fairly specific to the functions we associate them with, promising for manipulating deployed models with them.
Ablated Task C.F. Acc. W.T. Acc. V.T. Acc.
Country Facts 1.00 0.93 1.00
Word Translation 1.00 0.83 0.99
Verb Transformation 0.73 0.64 0.66
Table 5: Prediction accuracy over concept-relation pairs in each task when components from the same or another task are ablated. If the same task, cases where prompted concept-relation pair overlaps with ablated concept-relation pair are omitted from evaluation.
4 Related Work
Prior work has applied various methods to interpret large neural networks as compositions of subnetworks that perform specific functions (Olah et al., 2020; Elhage et al., 2021; Sharkey et al., 2025). Specialized circuits have been discovered for various functions of LLMs, including in-context learning (Olsson et al., 2022), numerical comparison (Hanna et al., 2023), indirect object recognition (Wang et al., 2023), and broader natural language fluency AlKhamissi et al. (2025). More recent work has developed automated circuit discovery methods to improve scalability (Goldowsky-Dill et al., 2023; Conmy et al., 2023; Hsu et al., 2025; Bhaskar et al., 2024). However, these approaches can be computationally expensive and yield circuits that are difficult to interpret.
As large neural networks like LLMs are thought to encode sparse features across polysemantic neurons (Elhage et al., 2022), dictionary learning techniques like SAEs promise greater interpretability by extracting monosemantic features from them (Bricken et al., 2023; Huben et al., 2024). This has allowed automated circuit discovery approaches to operate on these features instead of neurons (Marks et al., 2025), but high computational costs remain a barrier. Like our work, Li et al. (2025) used SAE feature coactivation patterns to analyze their geometry, finding spatial clustering of related concepts. Transcoders (Dunefsky et al., 2024) extend SAEs by learning to approximate layerwise computations. Cross-layer transcoders comprise the Anthropic circuit tracing tool (Ameisen et al., 2025), which visualizes LLM responses with sparse but still difficult to interpret computation graphs. Building on these efforts, we instead follow Li et al. to leverage SAE feature coactivation, applying node pruning based on activation density to automatically discover semantically coherent, context-consistent components that can influence LLMs’ relational reasoning outputs individually and in combination. Our approach offers an efficient framework for analyzing and controlling LLM behavior without exhaustive circuit tracing.
Orthogonally, factual knowledge is believed to reside in the feedforward layers of transformer-based LLMs. Geva et al. (2021) characterized these layers as key-value memories mapping textual patterns to vocabulary distributions. Dai et al. (2022) identified “knowledge neurons” whose activations correlate with specific facts. These insights informed approaches for editing factual knowledge stored in LLMs (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022). Geva et al. (2023) described the recall of factual associations as a three-step process of subject enrichment, relation propagation, and attribute extraction, while Hernandez et al. (2024) demonstrated that relation decoding in transformers can be approximated by simple linear transformations. Recent work on “knowledge circuits” has begun tracing causal pathways underlying factual recall (Yao et al., 2024; Ou et al., 2025). Most relevant to this work, Merullo et al. (2024) demonstrated that LLMs implement word2vec-style vector arithmetic through their feedforward layers to solve some relational tasks. Our work extends this effort by identifying finer-grained modular groups of features that represent human-interpretable concepts and relations which are similarly composed to form responses.
5 Conclusion
By tracing sparse feature coactivation patterns across layers, we uncover semantically coherent, context-consistent components that encode both concrete entities and concepts as well as abstract relations. We find that ablating or amplifying these components reliably alters model behavior. Moreover, composing country and relation components induces compound counterfactual responses. The layer-wise distribution of components reveals a hierarchical organization: entity features emerge in early layers, while relational features cluster in deeper layers. Overall, our method provides a lightweight, interpretable framework for analyzing and steering LLM behavior without full-scale circuit tracing. This work advances scalable interpretation and control of LLM behavior by leveraging pretrained SAEs to efficiently identify functional components. Given the increased adoption of LLMs by humans on a greater variety of tasks, such capabilities are vital for sustaining a human-machine enterprise that is not only reliable and safe, but also transparent, aligned with human goals, and resilient to misuse and misinterpretation.
Limitations
Some key limitations of our study warrant acknowledgment:
Limited datasets.
First, due to computational constraints, our investigation was constrained to only 3 multi-relational reasoning tasks with small numbers of examples. When it comes to the task space, we attempted to add some additional tasks around Spanish verb conjugation, factual information about Nobel Prize winners, and taxonomical information about animals, [^10] More discussion about animal taxonomy results in Appendix K. but Gemma 2 2B could not perform these tasks accurately off the shelf (for the former, this is likely due to Gemma not being trained on much Spanish data), thus we would not expect to identify semantic modules for these tasks. That said, the tasks we explored in this paper are diverse, representing factual and lexicosemantic knowledge, and similar forms of these tasks are also considered in prior relevant works Merullo et al. (2024); Todd et al. (2024); Hernandez et al. (2024). As such, we expect that our analytical framework generalizes to other domains.
When it comes to the datasets’ sizes, scaling our method up to large numbers of examples causes some complications which we intentionally chose to avoid. First, increasing the dataset sizes causes a combinatorial explosion for the reporting of steering results, which currently involve steering from each concept/relation/concept-relation pair to every other concept/relation/concept-relation pair. Consequently, with a larger dataset, any steering results reported would have to be based on subsets of such combinations, e.g., specific interchange interventions as done in the RAVEL dataset Wu et al. (2024). However, this has a tradeoff of providing a less clear view into the composition of the semantic modules identified in our work.
This is also nontrivial due to the requirement of manually selecting components based on various heuristics outlined in Section 3. While we believe it may be possible to automate these heuristics or even optimize component selection in future work, this would still take a significant amount of time and computational power. Nevertheless, we believe the examples in our datasets comprise sufficiently representative samples of the domains, e.g., countries around the world, animals from various classes, orders, and families, words of various syntactic categories, and verbs with both common and irregular past tense transformations.
Limited selection of interpretability methods and models.
Computational constraints also restricted our analysis to standard JumpReLU sparse autoencoders (Rajamanoharan et al., 2024b), which can suffer from inconsistent feature quality, poor reconstruction, and weak functional alignment. These limitations have motivated various improved architectures (Braun et al., 2024; Rajamanoharan et al., 2024a, b; Dunefsky et al., 2024), which may yield even more refined functional module identification under our approach. As a proof-of-concept, we extended our translation analysis to Gemma 2 2B transcoders (Dunefsky et al., 2024), successfully rediscovering some causal features for translation languages that Hanna et al. manually select, and achieving about 27% steering success rate with our extracted components. [^11] See Appendix L for more details. This provides promise that this approach can work with broader interpretability methods than SAEs.
Related to this, most of our results were limited to one model (Gemma 2 2B), except for the supplementary results on the country facts task with Gemma 2 9B in Appendix A. Our choice of LLMs was constrained to those with pretrained SAEs available via Neuronpedia. Further, we did not use smaller models like GPT-2, as they demonstrated a weaker grasp of the factual concepts under investigation (e.g., confusing Lagos, Nigeria’s largest city, with its capital) and an inability to follow in-context instructions to shorten their answers. As we continue to improve this work, we hope to expand the results to more models within the bounds of our computational constraints.
Focus on sparsely activating features.
Lastly, our analysis focused exclusively on features with low activation density. Intriguingly, preliminary experiments revealed that ablating high-density features produces syntactically and semantically incoherent outputs, suggesting these features serve critical yet unexplored computational functions that merit further investigation.
Appendix A Gemma 2 9B Results
We replicated country fact task ablation and steering using Gemma 2 9B. Results closely mirrored those observed with Gemma 2 2B. The model showed context consistency, with similar country components across relations (Figure 5) and similar relation components across countries (Figure 6). High success rates were achieved for country (93%), relation (97%), and composite (92%) steering (see Tables 68).
Figure 5: Gemma 2 9B China components extracted from capital and currency prompts.
Figure 6: Gemma 2 9B language components extracted from China and Nigeria prompts.
Country CN FR DE JP NG PL RU ES UK US Average
Success Rate 1.00 0.78 0.85 1.00 0.89 0.81 1.00 0.96 1.00 0.96 0.93
Table 6: Gemma 2 9B steering success rates for each target country across all 27 prompts that do not already query that country.
Relation Capital Currency Language Average
Success Rate 0.90 1.00 1.00 0.97
Table 7: Gemma 2 9B steering success rates for each target country fact across all 20 prompts that do not already query that country fact.
Ctry. / Rel. CN FR DE JP NG PL RU ES UK US Avg.
Capital 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Currency 1.00 1.00 1.00 0.72 1.00 0.50 0.11 0.94 1.00 1.00 0.83
Language 0.50 1.00 1.00 1.00 1.00 0.78 1.00 1.00 1.00 1.00 0.93
Average 0.83 1.00 1.00 0.91 1.00 0.76 0.70 0.98 1.00 1.00 0.92
Table 8: Gemma 2 9B composite steering success rates for each target country-fact pair across all 18 prompts that do not already query the target country or fact.
Appendix B Full Task Details
In this appendix, we provide the full list of concepts included in our analysis for each task. Additionally, as our accuracy and steering success evaluations importantly accounted for multiple possible correct answers to prompts, we provide lists of correct answers for each explored task’s concept-relation pairs below. These lists were refined through inspecting correct answers counted as incorrect in intermediate results. Importantly, the off-the-shelf tested models achieve 100% accuracy on all tasks under the few-shot prompts used to identify components, as well as the zero-shot prompts used for ablation and steering experiments. The model’s inconsistent performance may be another reason for the relatively low steering success rates on this task.
B.1 Country Facts
For the country facts, the full list of countries and their corresponding facts is below:
  • Capital:
    • China $\rightarrow$ Beijing
    • France $\rightarrow$ Paris
    • Germany $\rightarrow$ Berlin
    • Japan $\rightarrow$ Tokyo
    • Nigeria $\rightarrow$ Abuja
    • Poland $\rightarrow$ Warsaw
    • Russia $\rightarrow$ Moscow, Москва
    • Spain $\rightarrow$ Madrid
    • UK $\rightarrow$ London
    • USA $\rightarrow$ Washington
  • Currency:
    • China $\rightarrow$ Yuan, ¥, RMB, CNY, renmin
    • France $\rightarrow$ Euro, Franc, €
    • Germany $\rightarrow$ Euro, €, Mark
    • Japan $\rightarrow$ Yen, ¥
    • Nigeria $\rightarrow$ Naira, ₦
    • Poland $\rightarrow$ Złoty, Zloty, PLN, Zlo
    • Russia $\rightarrow$ Ruble, , RUB
    • Spain $\rightarrow$ Euro, €
    • UK $\rightarrow$ Pound, £
    • USA $\rightarrow$ Dollar, \$
  • Language:
    • China $\rightarrow$ Mandarin, Chinese
    • France $\rightarrow$ French, Français
    • Germany $\rightarrow$ German, Deutsch
    • Japan $\rightarrow$ Japanese, 日本語
    • Nigeria $\rightarrow$ English
    • Poland $\rightarrow$ Polish
    • Russia $\rightarrow$ Russian
    • Spain $\rightarrow$ Spanish
    • UK $\rightarrow$ English
    • USA $\rightarrow$ English
B.2 Word Translation
For translation, our evaluation scheme accounts for various aspects:
  • Parts of speech, e.g., poisson (fish) and pêcher (to fish) as correct French translations of fish)
  • Verb conjugations, e.g., amo and amas as correct Spanish translations of love (despite this verb taking the same form in English regardless of grammatical first or second person)
  • Semantic relevance, e.g., coureur (runner) as a valid French translation of run despite being syntactically incorrect
These choices allow us to separate syntactic errors from semantic errors (the latter of which we focus on). The full list of correct answers is below:
  • Spanish:
    • beautiful $\rightarrow$ hermosa, hermoso, bello, bella, bonito, bonita, lindo, linda, belleza, bellisima
    • cat $\rightarrow$ gato, gata, felino
    • dog $\rightarrow$ perro, perra
    • fish $\rightarrow$ pez, pescar, pesco, pescas, pescan, pescado, pesca, faenar
    • good $\rightarrow$ buen, bueno, buena, bien, buenos, buenas
    • here $\rightarrow$ aquí, acá, aqui, Aqui
    • learn $\rightarrow$ aprender, aprendo, aprendes, aprenden, aprende, Saber
    • love $\rightarrow$ amar, amo, amas, aman, encanta
    • red $\rightarrow$ rojo, roja
    • run $\rightarrow$ correr, corro, corres, corren, corrida, corre
    • you $\rightarrow$ tú, eres
  • French:
    • beautiful $\rightarrow$ belle, beau, bel
    • cat $\rightarrow$ chat, chatte, félin
    • dog $\rightarrow$ chien, chienne
    • fish $\rightarrow$ poisson, pêcher, pêche
    • good $\rightarrow$ bien, bon, grand
    • here $\rightarrow$ ici
    • learn $\rightarrow$ apprendre, apprends, apprenant, savoir
    • love $\rightarrow$ amour, aimer, aime, aiment, aimes, aimez, aim
    • red $\rightarrow$ rouge
    • run $\rightarrow$ courir, cours, course, couler, coureur, A pied
    • you $\rightarrow$ tu, toi, vous, on, te
  • German:
    • beautiful $\rightarrow$ schön
    • cat $\rightarrow$ Katze
    • dog $\rightarrow$ Hund
    • fish $\rightarrow$ Fisch
    • good $\rightarrow$ gut, gute, guten, Gut
    • here $\rightarrow$ hier
    • learn $\rightarrow$ lernen, lerne, lernst, lernt, lern
    • love $\rightarrow$ lieben, liebe, liebst, liebt, lieb
    • red $\rightarrow$ rot
    • run $\rightarrow$ laufen, laufe, läufst, läuft, lauf, rennen, renne, rennst, rennt, renn, Flucht
    • you $\rightarrow$ du, Sie
B.3 Verb Transformation
For verb transformation, the full list of verbs and their corresponding transformations is below:
  • Synonym:
    • break $\rightarrow$ shatter, destroy, split, pause, end, fracture, stop
    • focus $\rightarrow$ concentrate, center, attention, attend, sharpen
    • hide $\rightarrow$ conceal, camouflage, bury, stash, secret
    • include $\rightarrow$ incorporate, contain, comprise, cover, involve, add, insert, append, package
    • like $\rightarrow$ enjoy, similar, alike, love, prefer
    • possess $\rightarrow$ have, hold, own, control, dominate, influence, keep, succeed
    • sink $\rightarrow$ submerge, dip, drop, descend, plunge, drain, basin, fall, destination
    • understand $\rightarrow$ realize, realise, grasp, know, comprehend, get
  • Antonym:
    • break $\rightarrow$ fix, mend, repair, continue, resume, make, whole, hold
    • focus $\rightarrow$ distract, unfocus, ignore, blur, lose, disperse, diversion
    • hide $\rightarrow$ expose, reveal, flaunt, show
    • include $\rightarrow$ exclude, omit, remove, delete
    • like $\rightarrow$ dislike, hate, avoid
    • possess $\rightarrow$ dispossess, lose, want, need, lack, abandon, miss, disarray, fail, not have
    • sink $\rightarrow$ float, rise, ascend
    • understand $\rightarrow$ misunderstand, misinterpret, ignore, not understand
  • Past Tense:
    • break $\rightarrow$ broke, broken
    • focus $\rightarrow$ focused
    • hide $\rightarrow$ hid, hidden
    • include $\rightarrow$ included
    • like $\rightarrow$ liked
    • possess $\rightarrow$ possessed
    • sink $\rightarrow$ sank, sunk
    • understand $\rightarrow$ understood
  • Capitalize:
    • break $\rightarrow$ Break
    • focus $\rightarrow$ Focus
    • hide $\rightarrow$ Hide
    • include $\rightarrow$ Include
    • like $\rightarrow$ Like
    • possess $\rightarrow$ Possess
    • sink $\rightarrow$ Sink
    • understand $\rightarrow$ Understand
  • First Letter:
    • break $\rightarrow$ B
    • focus $\rightarrow$ F
    • hide $\rightarrow$ H
    • include $\rightarrow$ I
    • like $\rightarrow$ L
    • possess $\rightarrow$ P
    • sink $\rightarrow$ S
    • understand $\rightarrow$ U
Like translation, the synonym and antonym relations similarly account for a variety of interpretations of the input word, as well as some answers which have a different part of speech than the input word.
Appendix C Prompt Templates
In this appendix, we list component selection and steering evaluation prompt templates for all tasks’ relations.
C.1 Country Facts
Capital city.
For component selection, we use:
The capital city of Peru is Lima. The capital city of South Korea is Seoul. The capital city of Saudi Arabia is Riyadh. The capital city of {in-prompt country} is
For steering evaluation, we use:
Q: What is the capital city of {in-prompt country}? Answer directly (two words max).
A:
Currency.
For component selection, we use:
The currency of Peru is the Sol. The currency of South Korea is the Won. The currency of Saudi Arabia is the Riyal. The currency of {in-prompt country} is the
For steering evaluation, we use:
Q: What is the currency of {in-prompt country}? Answer directly (two words max).
A:
Language.
For component selection, we use:
The main language in Peru is Spanish. The main language in South Korea is Korean. The main language in Saudi Arabia is Arabic. The main language in {in-prompt country} is
For steering evaluation, we use:
Q: What is the main language in {in-prompt country}? Answer directly (two words max).
A:
C.2 Word Translation
Spanish.
For component selection, we use:
The Spanish word for “sing" is cantar. The Spanish word for "he" is él. The Spanish word for "bird" is pájaro. The Spanish word for "{in-prompt word}" is
For steering evaluation, we use:
Q: What is the Spanish word for "{in-prompt word}"? Answer directly (two words max).
A:
French.
For component selection, we use:
The French word for “sing" is chanter. The French word for "he" is il. The French word for "bird" is oiseau. The French word for "{in-prompt word}" is
For steering evaluation, we use:
Q: What is the French word for "{in-prompt word}"? Answer directly (two words max).
A:
German.
For component selection, we use:
The German word for “sing" is singen. The German word for "he" is er. The German word for "bird" is Vogel. The German word for "{in-prompt word}" is
For steering evaluation, we use:
Q: What is the German word for "{in-prompt word}"? Answer directly (two words max).
A:
C.3 Verb Transformation
Synonym.
For component selection, we use:
A synonym of throw is toss. A synonym of go is move. A synonym of find is discover. A synonym of {in-prompt verb} is
For steering evaluation, we use:
Q: What is a synonym of {in-prompt verb}? Answer directly (one word max).
A:
Antonym.
For component selection, we use:
An antonym of throw is catch. An antonym of go is stop. An antonym of find is lose. An antonym of {in-prompt verb} is
For steering evaluation, we use:
Q: What is an antonym of {in-prompt verb}? Answer directly (one word max).
A:
Past tense.
For component selection, we use:
The past tense of throw is threw. The past tense of go is went. The past tense of find is found. The past tense of {in-prompt verb} is
For steering evaluation, we use:
Q: What is the past tense of {in-prompt verb}? Answer directly (one word max).
A:
Capitalize.
For component selection, we use:
The word throw with the first letter capitalized is Throw. The word go with the first letter capitalized is Go. The word find with the first letter capitalized is Find. The word {in-prompt verb} with the first letter capitalized is
For steering evaluation, we use:
Q: Capitalize the word {in-prompt verb}. Answer directly (one word max).
A:
First letter.
For component selection, we use:
The first letter of throw is T. The first letter of go is G. The first letter of find is F. The first letter of {in-prompt verb} is
For steering evaluation, we use:
Q: What is the first letter of the word {in-prompt verb}? Answer directly (one word max).
A:
Appendix D Additional Component KL Divergence Plots
In Figure 2, we plotted all sparse components for love and the three languages in the word translation task. We provide some additional examples in this appendix. Specifically, Figure 7 visualizes the component KL divergences for China and the capital city, currency, and language relations in the country facts task. Further, Figure 8 visualizes the component KL divergences for hide and the capitalize, first letter, and past tense relations in the verb transformation task.
Figure 7: Plot of component feature counts versus KL divergence between pre- and post-ablation output token distributions for sparse components extracted for the capital city, currency, and language of China.
Figure 8: Plot of component feature counts versus KL divergence between pre- and post-ablation output token distributions for sparse components extracted for the capitalization, first letter, and past tense of hide.
Appendix E Country Facts Component Feature Description Visualizations
Figure 9 visualizes the feature descriptions for selected components in the country fact task.
Refer to captionChina Component
Refer to captionCapital Component
Refer to captionCurrency Component
Language Component
Figure 9: Word clouds for LLM-generated descriptions of SAE features within the China, capital, currency, and language components.
Appendix F Additional Concept and Relation Component Visualizations
In this appendix, we include some additional visualizations of components extracted from studied tasks. Figures 10 and 11 visualize selected components from the word translation task, while Figure 12 visualizes selected components from the verb transformation task.
Figure 10: Components representing Spanish, French, and German translation (from left to right).
Figure 11: Components representing the translated words beautiful, dog, and love (from left to right).
Figure 12: Components representing the capitalize (upper left) and first letter (upper right) transformations in the context of the verb like, and the components representing the verb like in the context of the capitalize (middle row) and first letter (bottom row) transformations.
Appendix G Steering Details and Hyperparameters
For individual concept and relation steering, we selected steering strengths $\alpha_{c},\alpha_{r}$ from $\{k\cdot 0.05:k\in\mathbb{Z}\}\cap(0,1]$ that achieved the highest respective success rates. For composite steering, we selected the $(\alpha_{c}^{\prime},\alpha_{r}^{\prime})$ pair from $\{\alpha_{c}-0.05,\alpha_{c},\alpha_{c}+0.05\}\times\{\alpha_{r}-0.05,\alpha_{r},\alpha_{r}+0.05\}$ that achieved the highest success rate. This procedure yielded the following parameters for each task:
  • Country facts: $\alpha_{c}=0.1$, $\alpha_{r}=0.45$, $\alpha^{\prime}_{c}=0.15$, and $\alpha^{\prime}_{r}=0.45$
  • Word translation: $\alpha_{c}=0.05$, $\alpha_{r}=0.45$, $\alpha^{\prime}_{c}=0.05$, and $\alpha^{\prime}_{r}=0.4$
  • Verb transformation: $\alpha_{c}=0.05$, $\alpha_{r}=0.45$, $\alpha^{\prime}_{c}=0.055$, and $\alpha^{\prime}_{r}=0.5$
Appendix H Supplementary Ablation and Steering Next Token Distributions
In this appendix, we provide additional detailed examples of how the next token distributions change under ablating and steering concept and relation components.
H.1 Individual Concept and Relation Ablation
Extending Table 1, we provide three comprehensive sets of examples for ablating individual concepts and relations: Table 9 (for country facts), Table 10 (for word translation), and Table 11 (for verb transformations). While the results largely mirror those highlighted in Table 1, we observe that when ablating the first letter transformation, the model still ranks tokens representing single letters highly, and many of those letters are other letters besides the first in the prompted verb (e.g., K, E, and A in the context of break). This suggests that the first letter component specifically captures first letters, and not other letters in prompted words; as such, ablating it still leaves a signal to the model to extract (non-first) letters.
Capital Currency Language
Original Ctry. Abl. Fact Abl. Original Ctry. Abl. Fact Abl. Original Ctry. Abl. Fact Abl.
Beijing 97. Madrid 39. the 12. Yuan 80. Euro 64. Yuan 59. Mandarin 59. Spanish 49. Chinese 24.
Be .38 Warsaw 10. China 6.7 Ren 14. Lira 20. Ren 14. Chinese 37. English 22. China 18.
Peking .35 Rome 9.5 Beijing 6.4 RMB 1.8 Krone 3.1 Yen 1.9 English .69 French 6.7 Mandarin 11.
Shanghai .23 Paris 6.1 Shanghai 5.0 Yen .98 Franc 2.4 P 1.7 also .43 German 4.0 also 3.5
Xi .11 Berlin 5.0 a 2.7 yuan .53 Peso 1.2 Ba 1.5 Put .22 Italian 2.7 a 2.9
Abuja 85. New 7.7 Lagos 17. Naira 93. Franc 16. Naira 62. English 72. English 50. Nigeria 40.
Lagos 11. Islamabad 7.3 the 12. N 2.9 Euro 9.1 Dollar 6.0 Ha 11. French 31. English 14.
.38 Kathmandu 6.0 Nigeria 7.1 Nai 2.5 D 7.4 Niger 4.3 Yoruba 4.7 Spanish 2.1 Nigerian 9.8
Nigeria .29 Delhi 5.7 called 5.1 naira .36 Pound 5.1 Currency 2.4 Igbo 2.4 Arabic 2.0 2.9
.27 Tehran 4.9 Abuja 3.4 K .23 Krone 4.8 Nai 1.4 Nigerian 1.6 Dutch 1.1 also 2.3
Table 9: Top 5 output tokens and their likelihoods for China (top) and Nigeria (bottom) across all three relations, before and after ablating the relevant country or relation component in Gemma 2 2B. Underscore prefixes of tokens are omitted for space.
Spanish French German
Original Word Abl. Lang. Abl. Original Word Abl. Lang. Abl. Original Word Abl. Lang. Abl.
Amor 41. " 5.9 Love 42. Amour 35. Le 15. Love 42. Liebe 66. Du 9.0 love 24.
amor 19. El 4.9 love 14. amour 26. La 6.4 I 11. Lie 4.8 " 6.6 Love 16.
Amo 7.0 La 4.4 I 13. A 5.7 " 6.1 love 10. Ich 4.4 Die 4.5 Liebe 11.
amo 5.1 Malo 3.2 " 5.3 J 3.8 le 3.7 " 5.8 " 3.4 das 4.2 I 9.6
" 2.9 la 2.5 LOVE 2.7 " 2.7 la 3.3 LOVE 2.5 Herz 1.7 Der 3.8 " 6.7
rojo 38. La 4.1 Red 29. Rouge 43. La 6.3 Red 38. Rot 51. " 5.2 red 30.
Rojo 34. " 3.5 red 16. rouge 31. " 6.1 red 20. rot 28. Gel 3.2 Red 18.
rojo 2.1 Bol 3.2 The 6.4 Rou 4.7 la 3.8 The 6.6 rote 5.8 Sch 2.5 The 6.1
Ro 2.0 la 2.8 " 6.4 ROU 2.6 Le 2.9 Rouge 5.8 Rote 3.5 1.9 " 5.8
Roja 1.9 2.7 4.1 Le 2.2 le 2.7 " 4.3 ro 2.0 die 1.5 2.9
Table 10: Top 5 output tokens and their likelihoods for love (top) and red (bottom) across all three languages, before and after ablating the relevant word or language component. Underscore prefixes of tokens are omitted for space.
Past Tense Capitalize 1st Letter
Original Verb Abl. Trans. Abl. Original Verb Abl. Trans. Abl. Original Verb Abl. Trans. Abl.
Broke 37. - 9.9 Past 8.7 Break 22. Capital 10. break 20. B 62. T 40. K 30.
broke 27. T 5.0 T 7.7 break 13. The 7.6 Break 7.1 b 17. C 11. E 14.
Broken 7.6 Was 4.6 ( 7.4 BREAK 8.3 4.2 5.1 A 1.9 A 10. k 13.
BRO 6.2 To 4.5 t 5.6 Capital 5.7 I 3.6 capital 4.3 E 1.5 O 9.9 A 5.7
Bre 2.8 O 4.3 E 5.1 The 5.0 Yes 2.3 The 3.3 Break 1.3 t 6.7 e 5.1
liked 48. was 8.7 like 8.4 Like 21. Capital 5.6 like 12. L 38. E 48. A 13.
Liked 34. have 7.6 Past 8.3 like 12. The 4.7 6.1 l 8.7 e 9.8 I 6.8
- 1.1 Have 7.0 T 8.2 Capital 8.6 Yes 4.5 capital 5.6 A 6.2 H 9.5 a 6.7
Loved 1.0 had 5.9 t 7.2 LIKE 4.3 No 4.1 the 4.6 I 4.1 F 6.6 E 4.6
Liked 1.0 - 5.9 Like 7.2 The 4.0 3.3 Like 3.8 " 3.2 T 3.8 " 4.2
Table 11: Top 5 output tokens and their likelihoods for break (top) and like (bottom) across three verb transformations, before and after ablating the relevant verb or transformation component. Underscore prefixes of tokens are omitted for space.
H.2 Individual Concept and Relation Steering
Extending Table 3, we provide three comprehensive sets of examples for steering individual concepts and relations: Table 12 (for country facts), Table 13 (for word translation), and Table 11 (for verb transformations). Observations from these extended examples largely resemble what we observed in Table 4. Related to what we observed in Table 11, we observe that steering toward capitalize is only partly successful at promoting the capitalized verb “Break,” while other letters in break besides the first letter (B) compete with it.
Capital, CN Currency, CN Language, CN
$\cdot$, NG Currency, $\cdot$ Language, $\cdot$ $\cdot$, NG Capital, $\cdot$ Language, $\cdot$ $\cdot$, NG Capital, $\cdot$ Currency, $\cdot$
_Abuja 87. _Yuan 15. _Mandarin 32. _Naira 75. _Beijing 98. _Chinese .35 _English 71. _Beijing 96. _Yuan 15.
_Nigeria 4.6 _yuan 12. _Chinese 27. _naira 8.3 _Be .75 _Mandarin .33 _Yoruba 5.6 _Be 1.5 _Ren 10.
_Lagos 3.9 _RMB 11. _English 25. _Nai 3.7 _BE .34 _English .28 _Ha 4.8 _Peking .83 _RMB 6.8
_- .58 _Ren 5.6 _Spanish 2.2 _Nigeria 3.3 _Peking .27 _Simplified .82 _Nigeria 4.4 Beijing .42 _yuan 6.3
_ .57 _China 5.4 _mandarin 2.1 _Nigerian 2.4 Beijing .25 _Spanish .79 _Igbo 3.4 _BE .30 _The 5.3
Table 12: Top 5 output tokens and likelihoods for prompts about China’s capital (columns 1–3), currency (columns 4–6), and language (columns 7–9), after ablating an in-prompt country or fact component (row 1 headings) and amplifying a target country (Nigeria) or fact component (row 2 headings). Highlighted tokens are or may begin correct answers for steered components.
Spanish, love French, love German, love
$\cdot$, red French, $\cdot$ German, $\cdot$ $\cdot$, red Spanish, $\cdot$ German, $\cdot$ $\cdot$, red Spanish, $\cdot$ French, $\cdot$
_rojo 43. _A 11. _Liebe 65. _rouge 44. _amor 41. _Liebe 60. _rot 53. _amor 45. _A 11.
_Rojo 8.7 _l 9.3 _die 2.5 _Rouge 30. _el 9.5 _die 4.0 _Rot 33. _el 11. _l 9.3
_Sang 5.7 _amour 7.9 _zu 2.3 _Sang 2.5 _Amor 8.5 _„ 3.3 _rote 3.6 _" 7.4 _amour 7.9
_roja 2.8 _" 7.3 _„ 2.2 _Rou 1.8 _" 6.3 _lieben 1.8 _Rote 2.0 _a 3.2 _" 7.3
_rojo 1.8 _ 5.2 _Ich 1.7 _ 1.2 _A 2.8 _der 1.4 _Blut 1.2 _la 2.6 _ 5.2
Table 13: Top 5 output tokens and likelihoods for prompts to translate the word love into Spanish (columns 1–3), French (columns 4–6), and German (columns 7–9), after ablating an in-prompt word or language component (row 1 headings) and amplifying a target word (red) or language component (row 2 headings). Highlighted tokens are or may begin correct answers for steered components.
Past Tense, break Capitalize, break 1st Letter, break
$\cdot$, like Capitalize, $\cdot$ 1st Letter, $\cdot$ $\cdot$, like Past Tense, $\cdot$ 1st Letter, $\cdot$ $\cdot$, like Past Tense, $\cdot$ Capitalize, $\cdot$
_liked 43. _Past 17. _B 24. _Like 17. $\otimes$ 36. _break 19. _L 54. $\otimes$ 35. _" 6.8
_Liked 43. _Has 7.8 _( 8.2 _I 7.0 _surla 16. _B 18. _l 11. _surla 16. _K 6.4
_L 1.6 _Kick 4.4 _Past 5.8 _like 6.9 AddTagHelper 14. _b 7.6 _A 4.0 AddTagHelper 15. _Break 6.4
_Like 1.5 _ 4.2 _E 3.0 _LIKE 6.7 _These 7.0 _Break 6.1 _I 2.4 _These 7.2 _E 6.0
_- 1.1 _Have 4.0 _T 2.9 _The 6.0 _nahilalakip 5.7 _ 3.5 _" 2.1 _nahilalakip 6.3 _B 5.6
Table 14: Top 5 output tokens and likelihoods for prompts to get the past tense (columns 1–3), capitalized form (columns 4–6), and first letter (columns 7–9) of the word break, after ablating an in-prompt verb or transformation component (row 1 headings) and amplifying a target verb (red) or transformation component (row 2 headings). Highlighted tokens are or may begin correct answers for steered components. “s” refers to Old English long S.
H.3 Composite Concept and Relation Steering
Extending Table 4, we provide two full sets of examples for steering concepts and relations in composition: Table 15 (for country facts) and Table 16 (for word translation). Observations from these extended examples largely resemble what we observed in Table 4.
Capital, China Currency, China Language, China
Currency, Nigeria Language, Nigeria Capital, Nigeria Language, Nigeria Capital, Nigeria Currency, Nigeria
_Naira 34. _English 71. _Abuja 70. _English 93. _Abuja 49. _Naira 40.
_Nigeria 28. _French 5.1 _Lagos 26. _French 1.7 _Lagos 46. _Nigeria 18.
_naira 13. _Yoruba 4.1 _ 2.3 _Yoruba 1.5 _ 1.7 _Dollar 6.2
_N 4.1 _Spanish 3.8 _Nigeria .41 _Spanish 1.1 _Nigeria .78 _naira 4.7
_ 3.2 _Igbo 3.0 _… .30 _Igbo .55 _… .36 _ 3.7
Table 15: Top 5 output tokens and their likelihoods for prompts about China’s capital (columns 1–2), currency (3–4), and language (5–6), after ablating both in-prompt components and amplifying a target country-fact pair. Highlighted tokens are or may begin correct answers for steered components.
Spanish, love French, love German, love
French, red German, red Spanish, red German, red Spanish, red French, red
_rouge 38. _rot 31. _rojo 38. _rot 26. _rojo 45. _rouge 31.
_Rouge 5.2 _Rot 17. _el 9.1 _Rot 14. _el 9.8 _le 11.
_la 3.0 _rote 9.9 _roja 5.0 _rote 7.4 _roja 4.2 _la 4.9
_Sang 3.0 _die 3.7 _El 3.4 _die 4.6 _" 3.7 _ 4.5
_le 2.6 _der 1.8 _la 3.2 _„ 2.7 _la 2.6 _" 3.7
Table 16: Top 5 output tokens and their likelihoods for prompts to translate love into Spanish (columns 1–2), French (3–4), and German (5–6), after ablating both in-prompt components and amplifying a target word-language pair. Highlighted tokens are or may begin correct answers for steered components.
Appendix I Steering Results by Concept
Table 2 summarized steering success rates by relation in each task. Here, Tables 17-19 break down concept steering success rate by concepts, while Tables 20-22 break down composite steering success rates by concept-relation pairs.
Country CN FR DE JP NG PL RU ES UK US Average
Success Rate 1.00 1.00 1.00 1.00 1.00 0.78 1.00 0.78 1.00 1.00 0.96
Table 17: Steering success rate for each target country across all 27 prompts for other countries.
Word cat dog fish you beautiful good red here learn love run Average
Success Rate 0.60 0.87 0.47 0.57 0.93 0.83 0.90 0.47 1.00 1.00 0.63 0.75
Table 18: Steering success rate for each target translated word across all 30 prompts for other words.
Verb break focus hide include like possess sink understand Average
Success Rate 0.66 0.34 0.69 0.63 0.83 0.06 0.20 0.46 0.48
Table 19: Steering success rate for each target verb across all 35 prompts for other verbs.
Ctry. / Fact CN FR DE JP NG PL RU ES UK US Avg.
Capital 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Currency 1.00 1.00 1.00 0.72 1.00 0.50 0.11 0.94 1.00 1.00 0.83
Language 0.50 1.00 1.00 1.00 1.00 0.78 1.00 1.00 1.00 1.00 0.93
Average 0.83 0.98 1.00 0.91 0.87 0.76 0.70 0.98 1.00 1.00 0.90
Table 20: Composite steering success rates for each target country-fact pair across all 18 prompts that do not already query the target country or fact.
Word / Lang. cat dog fish you beautiful good red here learn love run Average
Spanish 0.45 0.55 0.10 0.25 1.00 0.25 0.65 0.55 0.70 1.00 0.20 0.52
French 0.70 0.80 0.05 0.75 0.95 0.85 0.60 0.45 0.95 0.80 0.30 0.65
German 0.45 0.30 0.00 0.45 0.75 1.00 0.65 0.60 0.95 1.00 0.40 0.60
Average 0.53 0.55 0.05 0.48 0.90 0.70 0.63 0.53 0.87 0.93 0.30 0.59
Table 21: Composite steering success rates for each target word-language pair across all 20 prompts that do not already query the target word or language.
Verb / Trans. break focus hide include like possess sink understand Average
Synonym 0.00 0.00 0.00 0.00 0.14 0.00 0.00 0.00 0.02
Antonym 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Past Tense 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Capitalize 0.00 0.00 0.64 0.18 0.50 0.04 0.00 0.11 0.18
First Letter 0.93 0.36 0.93 0.86 1.00 1.00 1.00 1.00 0.88
Average 0.19 0.07 0.31 0.21 0.33 0.21 0.20 0.22 0.22
Table 22: Composite steering success rates for each target verb-transformation pair across all 10 prompts that do not already query the target verb or transformation.
Appendix J Supplementary Results on Component Organization
Extending Figure 4 from Section 3.3, we provide additional KL divergences by node and layer in Figure 13 (for translated words), Figure 14 (for target translation languages), Figure 15 (for verbs to transformer), and Figure 16 (for verb transformations).
Figure 13: KL divergence between pre- and post-ablation output token distributions for each node in the love and red components from the word translation task, plotted by layer. Linear regression lines plotted in red.
Figure 14: KL divergence between pre- and post-ablation output token distributions for each node in the translation target language components, plotted by layer. Linear regression lines plotted in red.
Figure 15: KL divergence between pre- and post-ablation output token distributions for each node in the break and like components from the verb transformation task, plotted by layer. Linear regression lines plotted in red.
Figure 16: KL divergence between pre- and post-ablation output token distributions for each node in the verb transformation components, plotted by layer. Linear regression lines plotted in red.
Appendix K Supplementary Results on Taxonomical Reasoning
Beyond the tasks in the paper, we additionally attempted to apply our method to a dataset for the taxonomic class, order, and family of 26 animals. However, we excluded these results from the paper after discovering that Gemma 2 2B could not perform this task perfectly off-the-shelf, thus steering based on extracted components also performed poorly.
Interestingly, however, causally impactful components for animal species exhibited a very strong hierarchical structure across the three taxonomy tasks. Generally, the component for a given species at the class level was a subgraph of the component for the same species at the order level, and the order level component a subgraph of the family level component. (17) This can be phrased as a general observation that as we move from more general to more specific categories, components increase in size by adding features.
We were able to reconstruct a reasonably accurate taxonomic hierarchy of the animals in our dataset by applying agglomerative clustering on the Jaccard distance matrix between family components of different species. The taxonomy is very accurate at class level, with some notable mistakes at order and family level. This coheres with previous results on LLMs’ general performance on taxonomy tasks Sun et al. (2024).
In future work, we intend to explore simpler tasks based on hierarchical knowledge.
Figure 17: Species components for Angelfish as extracted from class, order, and family tasks. Nodes present in all three are shown in blue, nodes present in order and family are shown in purple, and nodes present in just family are shown in red.
Appendix L Transcoder Analysis
Transcoders Dunefsky et al. (2024) are a major development in interpretability research. They operate by reconstructing the multilayer perceptron layer outputs from the inputs to the MLP layer.
To investigate whether our approach could be used to identify components for transcoders, we extend our translation task to the Gemma Scope 2 2B transcoders Lieberum et al. (2024) and focus on language components for French, German, and Spanish. The model’s outputs were quite robust to ablating top-activated features, so we were only able to identify causal components for dog, run, beautiful, learn, love, red, and good. In initial experiments, we found that for some prompts, we could ablate essentially all of the top-activated features without changing the model’s output, so we attribute this poor performance at least in part to relevant information being carried in the error term.
We found that a correlation threshold of 0.7 was more suitable for identifying causal components in transcoders, by testing decreasing values in increments of 0.5 and inspecting the results. We then ran our algorithm as described, with the exception of skipping the density pruning step. The lower threshold caused us to identify very large components with quite a bit of overlap. We assigned only the nodes unique to each language to the final components used for ablation and steering.
We conducted a relation steering experiment, ablating the component identified for the original language and steering the component identified for another. Testing integer values up to 20, we found that 13 produced the best accuracy. Our final accuracy was 27%.
This performance is not as good as our method achieves with SAEs, and comes with quite a few limitations. However, through manual inspection we found that the same causal features that Hanna et al. identify in their language intervention demo notebook are present in our identified components. While our method may need refining and adaptation to work at scale with transcoders, based on these results we feel confident that feature coactivation can help identify relevant conceptual components in transcoders.
Appendix M License Information
We provide available licenses and terms of use for key artifacts employed in this work, including relevant links:
We have verified that this work acts in accordance with all available licenses and terms of use.